I think I'm being asked to complete the square:
w(6w+5)=4

Answer by ankor@dixie-net.com(22740)

w(6w+5)=4
:
6w^2 + 5w + ___ = 4
:
We want the coefficient of w^2 to be 1, divide each term by 6
w^2 + (5/6)w + ___ = 4/6
:
Complete the square
w^2 + (5/6)w + (25/144) = 4/6 + (25/144): 1/2 of (5/6)^2 = 25/144
:
(w + 5/12)^2 = (96/144) + (25/144)
:
(w + 5/12)^2 = 121/144
:
w + (5/12) = +/- Sqrt(121/144); a perfect square!
:
w + (5/12) = +/- (11/12)
:
Two solutions:
w = -(5/12) + (11/12)
w = +(6/12)
w = +1/2
and
w = -(5/12) - (11/12)
w = -(16/12)
w = -(4/3)
:
Did this help?